Doakes Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on direct labor-hours. The company based its predetermined overhead rate for the current year on the following data:

Total direct labor-hours 60,000
Total fixed manufacturing overhead cost $ 378,000
Variable manufacturing overhead per direct labor-hour $ 2.20

Recently, Job M843 was completed with the following characteristics:

Number of units in the job 60
Total direct labor-hours 120
Direct materials $ 630
Direct labor cost $ 2,400

The unit product cost for Job M843 is closest to: (Round your intermediate calculations to 2 decimal places.)
Answer #1

Predetermined overhead = estimated total manufacturing overhead/established total Amount of the allocation base = (378000+(2.20*60000))/60000= 510000/60000=8.50 per direct labor hour

Total cost of JobM843= direct material+ direct labor+manufacturing overhead applied = 630+2400+(8.50*120)=4050

Unit product cost = total cost÷ number of units=4050÷60= $67.50